Show that for any real number x, if x2
is odd, then x is odd.
1
Expert's answer
2021-08-18T16:17:34-0400

If x is odd then x=2k+1.

"x^2=(2k+1)^2=4k^2+4k+1=2(2k^2+2k)+1=2n+1."

Thus, "x^2" is odd.